5/22-23: MZ release from PSL
5/20-24: Exchanging cryo valves.
5/27-6/2: Y arm mass lock. (not done yet)
5/29: RT PC maintenance -> go to safe mode for all VIS connected Dolphin (specially BS)
- 6/3-6/14: Installation of TMS for ETMY
- 6/8: Engineering run
5/22AM~: tuned off cryostat cooler, finesse measurement for Xarm
5/22AM~: tuned off shield cooler, finesse measurement for Xarm
5/2: Cooling start
5/23AM~: tuned off cryostat cooler, finesse measurement for Yarm
5/17~: tuned off ETMY cryostat cooler
- Long term:
- 4/22-5/31: Yarm commissioning
- 5/27-6/30: DRMI commissioning
- 7/1-8/9: DRFPMI commissioning
- 8/12-10/4: DRFPMI ASC
- 10/7-11/15: Post Commissioning
- 11/18-12/13: Noise hunting
- TMS installation this week.
- Wednesday noon, we will have a Run meeting.
Report of work
- TMS installation started at ETMY.
- Pressure of ITMX was recovered by changing cryo valve at duct cooler.
- Needs finesse measurement for X arm.
- PRX (PRM+ITMX) was Locked yesterday.
- Michelson was Locked using AS port first time.
- PRMI might be locked last night.
- Sweeping modulation frequency to minimize RF-AM conversion.
- Coil balance for 4 type A, and measuring TF.
- DGS: maintenance for installation whitening filter for iMC. iMC will be stopped.
- Dolphin issue will be fixed today.